### v1

# Augmenting LLMs for General Time Series Understanding and Prediction

Abstract
Time series data is fundamental to decision-making in many crucial domains including healthcare, finance, and environmental science. However, analyzing this data often requires incorporating unstructured contextual information, answering domain-specific questions, and generating natural language explanations – capabilities that traditional time series models lack due to their inability to process text. While Large Language Models (LLMs) excel at contextual reasoning and knowledge integration, they struggle with numerical time series due to inefficient text-based representations and limited exposure to temporal data during pretraining. We address this gap by augmenting an LLM with specialized time series perception through a patch-based encoder-decoder architecture. We train this Time Series-augmented LLM (TsLLM) on a large corpus of over 2 million interleaved time series and text examples spanning diverse analysis tasks: forecasting with contextual information, time series question-answering, pattern explanation, classification with natural language outputs, and report generation. This training enables TsLLM to leverage both its language understanding and newly acquired temporal reasoning capabilities. While not designed to surpass specialized models on traditional benchmarks, TsLLM demonstrates strong performance on tasks requiring the integration of time series analysis with natural language – capabilities that existing approaches cannot provide. Our work establishes a new paradigm for time series analysis that bridges numerical computation and natural language understanding, democratizing access to sophisticated temporal reasoning through natural language interaction.
1 Introduction
Many real-world time series workflows require more than numerical prediction. Clinicians ask questions about electrocardiograms in the context of symptoms and medications, analysts forecast while conditioning on news and product feedback, and scientists need narrative explanations of temporal patterns and anomalies. Traditional time series models are effective for narrowly defined numerical tasks, but they operate exclusively on structured inputs and produce only numerical outputs. As a result, practitioners often rely on brittle pipelines or expert intermediaries to bridge the gap between temporal signals, unstructured context, and high-level analysis. Furthermore, these models require data science expertise to build and interpret, presenting a significant barrier to entry for domain experts. Across healthcare, finance, scientific research, and industrial monitoring, domain experts need systems that can reason about time series data through natural language, answer questions about temporal patterns, and provide explanations for their analyses. The rise of multimodal machine learning models has transformed how we interact with visual and audio data, yet time series remains isolated from these advances, despite being one of the most pervasive data modalities in scientific and medical settings.
Large language models (LLMs) possess the capabilities needed for these integrated tasks: contextual reasoning, few-shot learning from examples, domain knowledge integration, and natural language interaction. As a result, utilizing LLMs to perform certain time series analysis tasks holds considerable promise (Jin et al., 2024; Zhang et al., 2024). However, their weak abilities to perceive numerical time series data prevent LLMs from effectively applying their capabilities to temporal analysis (Tan et al., 2024; Merrill et al., 2024). Two structural issues explain this gap. First, text tokenization represents each numeric value with multiple tokens, inflating sequence length and obscuring temporal structure; this limits the amount of signal and context that fits in context windows and hinders pattern recognition. Second, our analysis of LLM pretraining corpora reveals they contain little high-quality time series data, so LLMs lack the exposure needed for temporal analysis.
To address these limitations, we introduce TsLLM, a large language model augmented with dedicated time series perception that enables unified, interleaved text-time series reasoning and generation. TsLLM adds a high-fidelity, patch-based encoder-decoder pretrained with a $\beta$-VAE objective to produce compact, information-dense representations of temporal signals, along with a scale-aware encoding that decouples shape from magnitude so the model can reason about absolute scale while maintaining numerical stability. Lightweight MLP adapters map these time series latents into the LLM’s embedding space, and special tokens allow seamless fusion: inputs and outputs can interleave text and time series segments, and the model treats all tasks as next-token prediction. This preserves the pretrained LLM’s language ability while adding precise temporal perception. TsLLM’s unified formulation of language and time series modeling enables functions that traditional time series models and text-only LLMs generally do not provide: contextual forecasting conditioned on unstructured information, question answering over numerical time series data, and natural language guided time series synthesis, as well as standard time series tasks: forecasting, classification, anomaly detection and more, without using task-specific heads. Few-shot in-context prompting with interleaved examples further allows the model to adapt to new datasets and tasks without additional fine-tuning.
Empirically, TsLLM demonstrates strong performance on tasks that require integrating temporal signals with language. TsLLM significantly outperforms frontier LLMs on time series question-answering (QA) benchmarks, matches or exceeds specialized systems for QA on electrocardiogram data, and beats both LLM-based methods and traditional models on a contextual forecasting benchmark. It also retains most of its performance on standard text-only LLM evaluations. On traditional time series benchmarks, TsLLM is competitive with dedicated models on forecasting (e.g., GIFT-Eval) and classification (e.g., UCR), even though it is not optimized for purely numerical tasks. Ablations show that few-shot prompting provides significant gains in most settings, and that the performance of TsLLM scales with the size of the pretrained LLM it is built on, suggesting that its performance can be improved by scaling up further. The primary contributions of this work are:
  1. 1.
    The TsLLM model architecture, which integrates a time series encoder and decoder with a pretrained LLM, and allows for interleaved text and time series inputs and outputs.
  2. 2.
    A high-fidelity patch-based encoder-decoder for time series signals that is pretrained on a diverse set of time series data with a $\beta$-VAE objective to produce compact, information-dense representations for use in the LLM.
  3. 3.
    A multi-stage training procedure that aligns the LLM with the encoder and decoder models, then trains it to solve a wide range of time series analysis tasks on diverse data, adding temporal perception abilities to the LLM while preserving its language competence.
  4. 4.
    Strong empirical results on benchmarks that require integrating temporal signals with language, as well as traditional time series tasks.
  5. 5.
    A large corpus of multimodal interleaved text and numerical time series sequences, comprising over two million examples and 20 billion tokens in total, which were collected, augmented, and synthesized for training TsLLM.
  6. 6.
    We are releasing both the data and code for TsLLM to enable further research in this area (links to come).
Our goal is not to replace specialized time series models on pure numerical tasks, as they are effective and efficient for time series-specific tasks. Rather, TsLLM targets problems that require integrating temporal analysis with domain knowledge and natural-language interaction. We view TsLLM as complementary to dedicated time series models and foundation models: it expands what practitioners can do with temporal data by enabling direct, language-based interaction that is grounded in the underlying signals.
2 Related Works
Multimodal LLMs
Multimodal Large Language Models (MLLMs) extend LLMs’ capabilities beyond text to process diverse data types (Yin et al., 2024). These models integrate modality-specific encoders (e.g., vision encoders like CLIP (Radford et al., 2021) or audio encoders like CLAP (Elizalde et al., 2023)), with LLM backbones (e.g. LLaMa (Touvron et al., 2023; Grattafiori et al., 2024) and Qwen (Wang et al., 2024)). Connector modules translate encoded features into the LLM’s input space, evolving from simple MLP projections in LLaVA (Liu et al., 2023, 2024a) to more sophisticated mechanisms employing learnable query tokens (Li et al., 2023) and added cross-attention layers (Alayrac et al., 2022). Training typically follows a multi-stage approach: alignment pre-training to bridge encoders with LLMs using large-scale paired datasets, followed by multimodal instruction tuning (Liu et al., 2024a; Dai et al., 2023) to teach the models to better understand natural language instructions from users and solve tasks. The field continues to expand, incorporating 3D point clouds (Xu et al., 2024), video (Zhang et al., 2023), and any-to-any modal generation (Wu et al., 2024). Despite this progress, effective integration of continuous numerical data, particularly time series, remains a challenge due to the lack of high-quality paired time series and text datasets, and the complex, highly contextual nature of time series signals. TsLLM aims to address this gap by building on the foundations of MLLM research and adapting it to time series analysis.
Time Series Foundation Models
Inspired by recent successes in natural language processing (Zhao et al., 2023) and computer vision (Awais et al., 2025), research in time series analysis is increasingly focused on developing dedicated Time Series Foundation Models (TSFMs). These models are pretrained on large, diverse numerical time series corpora to learn universal temporal representations, and apply them to solving a task. The transformer architecture is a common backbone, with “patching” being a key adaptation for efficiency and local feature extraction, as demonstrated by PatchTST (Nie et al., 2023) for forecasting. Architectural variants include encoder-only models like Moirai (Liu et al., 2024b) for universal forecasting, and decoder-only models such as TimesFM (Das et al., 2024) pretrained on billions of data points for strong zero-shot forecasting.
While the aim of this body of work is to develop foundation models for time series, the vast majority of the current literature focuses solely on forecasting. There are a few notable exceptions, namely UniTS (Gao et al., 2024), which aims for broader multi-task and multi-domain generalization on numerical data through a unified token representation that encodes the task specifications, and MOMENT (Goswami et al., 2024), which trains a backbone model along with lightweight decoders for multiple tasks. Despite significant advances in learning from a broad set of numerical time series, these TSFMs lack the world knowledge and language processing capabilities that LLMs have to interpret rich, unstructured textual context, which is TsLLM’s goal.
Time Series LLMs
Distinct from efforts in dedicated TSFMs, applications of pretrained LLMs to time series analysis are rapidly evolving. The core objective is to leverage the rich contextual understanding of LLMs while adapting text-optimized architectures for continuous numerical data (Zhang et al., 2024). While early work demonstrated feasibility by prompting LLMs with text-based time series for zero-shot forecasting (Gruver et al., 2023), limitations in numerical precision and efficiency (Spathis and Kawsar, 2024) spurred the development of more sophisticated representation strategies. These include discretizing continuous series via quantization (Ansari et al., 2024; Zhang et al., 2025) or vocabulary expansion (Wang et al., 2025; Xie et al., 2024) (although at the cost of potential information loss (Chatzigeorgakidis et al., 2024)), and structuring the input through patching or multi-scale decomposition (Zhou et al., 2023; Jin et al., 2023; Chan et al., 2024; Kowsher et al., 2024). However, these methods often process time series and text as distinct, sequential blocks, focusing primarily on forecasting.
Recent work addresses input representation limitations by enforcing structural consistency within the LLM’s latent space to bridge modality gaps (Hu et al., 2025; Liu et al., 2025). Other approaches unlock reasoning potential through conversational and multi-agent systems for enhanced performance and interpretability (Xie et al., 2024; Lee et al., 2025). While these methods advance multimodal integration, they typically treat time series as a static input to be reasoned about, generating textual summaries or predictions as the final output. TsLLM addresses these critical limitations by interleaving time series and text within the inputs and, crucially, the outputs of a pretrained LLM. This enables deep connections between modalities and provides new flexibility in how problems can be constructed and solved.
3 Methodology
We propose TsLLM, a multimodal generative model that augments a pretrained decoder-only LLM with time series perception and generation abilities. Our design goals are: (i) to represent time series with high information density while preserving temporal structure and numeric fidelity; (ii) to enable interleaved text-time series inputs and outputs in a single autoregressive model; and (iii) to train the resulting model with a curriculum that first aligns modalities and then teaches a broad set of time series tasks under natural language instructions. TsLLM consists of four main components: (1) a patch-based time series encoder-decoder, (2) a pretrained LLM, (3) a pair of fusion layers that inject time series embeddings into the LLM token stream and maps LLM hidden states back to time series patches, and (4) a multi-stage training pipeline, as illustrated in Figure 1.
Our unified framework reformulates all time series tasks, including forecasting, classification, anomaly detection, and general question-answering, as autoregressive generation problems over interleaved sequences of text and time series tokens. This approach enables few-shot generalization to new domains through natural language instructions and in-context examples, while supporting rich contextual information such as domain knowledge, metadata, and even structured tabular data. The following sections detail each component that are primary contributions of this work, namely, our patch-based time series encoding mechanism (Sec. 3.1), the architectural modifications for multimodal integration (Sec. 3.2), and our multi-stage training strategy (Sec. 3.3).
Figure 1: TsLLM architecture enabling novel interleaving of text (blue) and time series (green) inputs and outputs. The time series encoder-decoder has been pretrained to produce high-quality time series patch representations, which our multi-stage alignment process (snowflake = frozen, fire = trained) bridges with the LLM to enable unified text and time series processing.
3.1 Time Series Encoding and Decoding
TsLLM depends on high-quality representations of time series signals for strong performance. However, current LLMs split numbers into multiple small tokens, making it hard to capture numerical relationships and resulting in inefficient, lengthy text-based sequences, with each number using four to eight tokens including delimiters (Spathis and Kawsar, 2023). This limits the amount of contextual information, e.g., covariates and few-shot examples, that can fit in the context window, and requires an alternative data encoding.
Patch-Based Encoding Strategy
We solve this representation challenge through a patch-based encoding mechanism that preserves temporal structure while achieving dramatic compression. Our approach segments time series into fixed-length patches of $p$ consecutive points, transforming each patch into a single dense representation. This design decision balances three competing objectives: (1) capturing sufficient local temporal patterns within each patch, (2) maintaining computational efficiency through reasonable sequence lengths, and (3) preserving fine-grained information necessary for precise analysis. The encoding process operates in two stages. First, each patch $\mathbf{x}_{i:i+p}$ undergoes local feature extraction through a learned MLP: $\mathbf{z}_{i}^{\text{local}}=\text{MLP}_{\text{patch}}(\mathbf{x}_{i:i+p})$, where the MLP captures patterns like trends, oscillations, and anomalies within the patch window. Second, these local representations are contextualized through a decoder-only transformer that models dependencies across patches: $\mathbf{z}_{i}=\text{TransformerEncoder}(\mathbf{z}_{1}^{\text{local}},...,\mathbf{z}_{n}^{\text{local}})_{i}$. The causal attention mechanism ensures compatibility with autoregressive generation, enabling the model to generate time series continuations naturally.
Architecture Details
Our encoder-decoder architecture employs several design choices optimized for time series representation. We use learned positional encodings rather than sinusoidal variants, allowing the model to adapt to different sampling frequencies and temporal scales. Pre-normalization and GELU activations improve training stability when handling diverse signal magnitudes. The patch size $p=32$ was selected through extensive ablation studies, providing an optimal trade-off: smaller patches preserve more detail but increase sequence length, while larger patches risk losing important local features.
For multivariate time series, we adopt a channel-independent encoding strategy where each univariate component is processed separately. This design choice, while seemingly restrictive, offers significant advantages: it naturally handles varying numbers of channels, enables transfer learning across different sensor configurations, and simplifies the alignment between time series and text modalities. The LLM backbone subsequently learns cross-channel relationships through its attention mechanism, leveraging its pre-trained capabilities for modeling complex dependencies.
Real-world time series exhibit enormous variability in scale – from microvolts in neural recordings to millions in financial transactions. To ensure stable training across this range, we apply sequence-level z-score normalization before encoding: $\hat{\mathbf{x}}=(\mathbf{x}-\mu_{\mathbf{x}})/\sigma_{\mathbf{x}}$, where $\mu_{\mathbf{x}}$ and $\sigma_{\mathbf{x}}$ are computed over the entire input sequence. We pretrain the encoder-decoder pair using a $\beta$-VAE objective that balances reconstruction fidelity with representation quality. The loss function combines two terms:
$\mathcal{L}_{\text{VAE}}=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}[\log p_{\theta}(\mathbf{x}|\mathbf{z})]-\beta\cdot\text{KL}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))$
where the reconstruction term (MSE loss) ensures accurate signal recovery, while the KL divergence term regularizes the latent space to be smooth and well-structured. We employ a sigmoid-based schedule for $\beta$, starting near zero to prioritize reconstruction accuracy, then gradually increasing to encourage disentangled representations. This scheduling strategy prevents the common “posterior collapse” problem in VAE training while ensuring high-quality reconstructions. The encoder-decoder is pretrained on 50 million sequences from the GIFT-Eval Pretrain dataset (Aksu et al., 2024), comprising 230 billion time series points across 88 diverse domains, including healthcare, finance, and industrial sensors. This large-scale pretraining ensures our representations generalize across domains, temporal resolutions, noise levels, and time horizons.
3.2 TsLLM Architecture
Having established how we encode time series into dense representations, we now describe how these representations are integrated with a pretrained language model to create a unified multimodal system. Our architectural design faces a fundamental challenge: how to combine time series understanding with the rich language capabilities of modern LLMs without degrading either modality’s performance.
Foundation Model Integration
Rather than training a multimodal model from scratch, we build upon existing pretrained LLMs for both practical and technical reasons. First, the scarcity of high-quality time series data on the internet—our analysis shows less than 0.1% of web text contains meaningful temporal sequences—makes joint pretraining economically infeasible. Second, pretrained LLMs already possess sophisticated reasoning capabilities, world knowledge, and instruction-following abilities that would be difficult to replicate. Our architecture therefore, focuses on augmenting these capabilities rather than replacing them.
We designed TsLLM to be model-agnostic, compatible with any decoder-only transformer architecture. In our experiments (Sect. 4), we demonstrate strong results with Qwen2.5 (Wang et al., 2024), Llama 3.2 (Grattafiori et al., 2024), and SmolLM2 (Allal et al., 2024), showing that our approach generalizes across different model families and scales, and that TsLLM can leverage future improvements in base LLM capabilities to enhance its performance.
Adapter Layers
The adapter layers serve as the critical bridge between our time series representations and the LLM’s embedding space. These components must solve a challenging alignment problem: mapping from the encoder’s learned representations (optimized for reconstruction) to the LLM’s token embeddings (optimized for next-token prediction). Our adapter architecture employs a two-layer MLP with residual connections:
$\mathbf{h}_{\text{adapted}}=\mathbf{z}+\text{MLP}_{2}(\text{GELU}(\text{MLP}_{1}(\mathbf{z})))$
where $\mathbf{z}$ is the encoded time series representation. The residual connection preserves the original temporal information while the MLP learns the necessary transformations for LLM compatibility. During training, we keep the encoder frozen and only update the adapter weights, preventing catastrophic forgetting of the encoder’s learned representations. We experimented with more sophisticated architectures, including cross-attention mechanisms (where time series representations query LLM embeddings) and Perceiver-based adapters (uses learned queries to extract information). However, ablation studies showed that simple MLPs performed comparably with 3x fewer parameters and 2x faster inference. This efficiency is crucial for practical deployment, especially for long time series sequences.
Scale-Shape Decomposition
Time series in real-world applications span vastly different scales, e.g., heartbeats measured in millivolts, stock prices in thousands of dollars, and particle counts in millions. This variability presents unique challenges for multimodal models. Unlike traditional time series models that can normalize inputs and rescale outputs using known statistics, TsLLM must: (1) understand scale-dependent semantics (e.g., a heart rate of 200 bpm indicates tachycardia), and (2) generate time series at appropriate scales from text-only prompts (e.g., “Generate normal ECG signals”). Our solution decouples each time series into shape (normalized patterns) and scale (magnitude information) components. While the encoder processes normalized sequences $\hat{\mathbf{x}}=(\mathbf{x}-\mu_{\mathbf{x}})/\sigma_{\mathbf{x}}$ to extract shape features, we separately encode the normalization parameters $\mu_{\mathbf{x}}$ and $\sigma_{\mathbf{x}}$ to preserve scale information.
Model Fusion
Integrating time series representations into the LLM requires careful handling of the token vocabulary and embedding spaces. We extend the LLM’s vocabulary with a special “<|ts|>” token that acts as a placeholder for time series patches. During preprocessing, we identify time series segments in the input and replace each patch with a “<|ts|>” token, maintaining the sequential structure of the data. At the embedding layer, we implement a routing mechanism that replaces “<|ts|>” token embeddings with their corresponding time series representations:
$\mathbf{e}_{i}=\begin{cases}\text{TokenEmbed}(t_{i})&\text{if }t_{i}\neq\texttt{<|ts|>}\\ \mathbf{h}_{\text{adapted},j}+\mathbf{s}_{j}&\text{if }t_{i}=\texttt{<|ts|>}\end{cases}$
where $\mathbf{h}_{\text{adapted},j}$ is the $j$-th adapted time series representation and $\mathbf{s}_{j}$ is its corresponding scale embedding. This additive combination allows the model to modulate time series patterns based on their scale context. For generation, we implement bidirectional conversion: the LLM predicts “<|ts|>” tokens where time series should be generated, and we extract the corresponding hidden states for decoding. This design enables truly interleaved generation—the model can freely mix text and time series in its outputs based on the task requirements.
3.3 Training TsLLM
In this section, we describe how TsLLM is trained to handle the full spectrum of time series analysis tasks. Our training strategy must address several challenges: unifying diverse task formats, balancing modality-specific losses, and ensuring stable multi-stage optimization.
Figure 2: Illustrative examples of our unified framework where time series data is integrated with the text in both the inputs and outputs. This figure visualizes these embedded time series segments as plots for illustrative clarity. Examples demonstrate versatility across tasks like: (Top) contextual forecasting (e.g., ECG signals), (Middle) classification (e.g., sensor data), and (Bottom) report generation (e.g., COVID mobility).
Unified Task Formulation
A key feature of TsLLM is reformulating all time series tasks as autoregressive token generation problems. This unification enables knowledge transfer across tasks and simplifies the training process. Forecasting becomes natural continuation: given historical data as time series tokens, the model generates future time series tokens. The autoregressive nature means every time series in our training data implicitly teaches forecasting. Classification transforms into constrained text generation. Given a time series and the prompt “What type of activity is this?”, the model generates class labels as text. We use constrained decoding during inference to ensure outputs match valid class names. Anomaly detection employs two complementary approaches: (1) direct detection where the model outputs anomaly locations as text (eg. “Anomalies detected at timestamps 145-162, 203-210”), and (2) reconstruction-based detection where the model generates expected normal behavior, with deviations indicating anomalies.
Other tasks including segmentation, point-level classification, and imputation can similarly be formulated as text generation tasks. This unified formulation is powerful but introduces a challenge: dataset-specific semantics. A “normal” pattern in one domain may be anomalous in another. We address this through few-shot prompting, including domain examples in the context to provide task-specific guidance without fine-tuning.
Multi-Objective Training
Our training objective balances three components: $\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{\text{CE}}+\lambda_{2}\mathcal{L}_{\text{MSE}}+\lambda_{3}\mathcal{L}_{\text{scale}}.$ The cross-entropy loss $\mathcal{L}_{\text{CE}}$ applies to text tokens, maintaining language modeling capabilities. The MSE loss $\mathcal{L}_{\text{MSE}}$ supervises time series reconstruction through the decoder. The scale loss $\mathcal{L}_{\text{scale}}$ ensures accurate recovery of normalization parameters. We use adaptive weighting with $\lambda_{1}=1.0$, $\lambda_{2}=0.5$, and $\lambda_{3}=0.1$, determined through extensive hyperparameter search to balance modalities without one dominating.
Staged Training Strategy
The training process for TsLLM employs a carefully orchestrated three-stage approach, with each stage serving a specific purpose in building the model’s capabilities. Stage 1: Alignment (1B tokens) – We begin by training only the adapter and scale encoding layers while keeping both the encoder-decoder and LLM frozen. This stage uses raw time series sequences and simple synthetic tasks (estimating the mean, translating to and from text-based time series encodings, etc.) to establish the connection between modalities. By limiting the trainable parameters, we ensure stable alignment without disrupting pretrained representations. Stage 2: Continued Pretraining (8B tokens) – With alignment established, we unfreeze all parameters for comprehensive multimodal training. This stage uses our full training dataset, with conversational prompt-response data undersampled and reformatted as unstructured sequences. Stage 3: Supervised Fine-Tuning (20B tokens) – The final stage shifts to a conversational format, training on instruction-response pairs with loss computed only on responses. This stage transforms TsLLM from a base model for next-token prediction into an instruction-following model that can solve time series analysis tasks specified in natural language. This staged approach proved crucial for stability as early experiments with all parameters unfrozen from the start tended to have loss spikes, and performance plateaued early.
3.4 Data Curation and Synthesis
Training TsLLM to bridge time series and language modalities requires comprehensive data coverage across three categories: pure time series for encoder pretraining, paired time series with text for multimodal alignment, and synthetic tasks incorporating time series with natural language instructions and context for capability development. Our training corpus comprises 25 billion tokens across 2 million examples, carefully curated to address the scarcity of high-quality time series-text pairs while ensuring broad domain coverage. The data collection process addresses a fundamental challenge: while web-scraped text corpora contain billions of documents, less than 0.1% include meaningful temporal sequences, necessitating alternative strategies for collecting suitable data.
Existing Datasets
We leverage established benchmarks to provide foundational capabilities across diverse time series tasks. For large-scale pretraining, we utilize GIFT-Eval (Aksu et al., 2024) with 230 billion time series points spanning 88 domains including healthcare, finance, energy, and environmental monitoring. This scale ensures our encoder-decoder learns robust temporal representations across different sampling frequencies (from seconds to years) and sequence characteristics. Classification capabilities are developed using the UCR (Dau et al., 2018) and UEA (Bagnall et al., 2018) archives, together covering 158 datasets ranging from physiological signals like electrocardiograms to industrial sensor readings and human activity recognition.
For paired time series-text data, we incorporate specialized datasets that naturally combine both modalities. ECG-QA (Oh et al., 2023) provides 200,000 cardiac signal question-answering pairs spanning diagnostic queries, rhythm identification, and clinical interpretation tasks. The TextTime Corpus (Kim et al., 2024) contributes weather measurements paired with meteorological discussions and clinical notes synchronized with patient vital signs from intensive care units. These existing paired datasets, while valuable, represent only a small fraction of our training data, highlighting the need for augmentation strategies.
Dataset Augmentation
To expand beyond limited existing paired datasets, we develop novel augmentation strategies that create meaningful text-time series associations while preserving signal fidelity. Our financial dataset pairs 5 years of S&P 500 price histories with 62,000 contemporaneous news articles from financial news sources, temporally aligned to enable contextual forecasting where predictions must integrate both numerical trends and market events. Each example combines a 32-day price history with up to 4 full articles or 16 headlines, ensuring textual context precedes prediction windows.
For medical applications, we augment the PTB-XL ECG dataset (Wagner et al., 2022) containing 21,799 clinical recordings to create three specialized variants through controlled text generation. The contextual forecasting variant pairs each ECG with multiple clinical reports using templates ranging from basic summaries to detailed interpretative analyses. The classification variant combines signals with demographic narratives and symptom descriptions while excluding pre-computed features, forcing reliance on raw signal patterns. The report generation variant produces structured findings, narrative clinical stories, and differential diagnoses from ECG patterns. This augmentation strategy transforms a single dataset into 65,000 multimodal training examples covering diverse analytical perspectives.
Synthetic Data Generation
The scarcity of naturally occurring paired time series-text data necessitates large-scale synthesis to achieve comprehensive task coverage. We generate 2 million question-answer pairs (10 billion tokens) through systematic combination of time series samples with templated analytical tasks. Our synthesis pipeline samples time series from diverse domains, constructs task-specific prompts spanning trend analysis, anomaly detection, periodicity identification, statistical reasoning, and causal interpretation, then generates responses that provide detailed analytical narratives.
Beyond question-answering, we create foundational tasks that establish core capabilities: bidirectional conversion between numerical arrays and textual representations (testing precise value preservation), statistical computations ranging from simple aggregations to complex distributional analyses, and multi-scale temporal reasoning from local patterns to long-term trends. This synthetic data follows recent findings (Bansal et al., 2024) suggesting that larger quantities of diverse examples, even with occasional inaccuracies, provide more robust training signals than smaller curated datasets. Our synthesis process ensures coverage of edge cases, rare patterns, and complex reasoning chains that might be underrepresented in natural data.
Data Formatting and Task Unification
We unify diverse time series tasks through consistent prompt-response formatting that enables seamless interleaving of text and time series. As illustrated in Figure 2, each task follows a structured template combining natural language instructions, optional context, time series signals (represented as <|ts|> tokens), and expected outputs. For contextual forecasting, prompts include historical signals, metadata, and relevant textual context, with responses generating future time series continuations. Classification tasks provide dataset descriptions and few-shot examples, enabling in-context learning without task-specific fine-tuning. Report generation prompts combine time series with metadata, producing detailed textual analyses that describe patterns, identify anomalies, and provide domain-specific interpretations. This unified formatting allows TsLLM to treat all tasks as variations of autoregressive generation, leveraging its language modeling capabilities while maintaining precise numerical fidelity. Complete details on dataset composition, synthesis procedures, and formatting specifications are provided in Appendix C.
4 Results
To demonstrate TsLLM’s unique capabilities at the intersection of time series analysis and natural language understanding, we evaluate it across diverse tasks ranging from those requiring language integration (question-answering, contextual forecasting) to traditional time series benchmarks (classification, anomaly detection). This evaluation strategy reflects our aims: while we show competitive performance on traditional tasks, our primary focus is demonstrating capabilities that existing approaches cannot provide. We prioritize evaluation on time series question-answering, where TsLLM’s ability to combine temporal reasoning with domain knowledge is most evident. We also assess performance on traditional benchmarks to establish that augmenting an LLM with time series perception maintains competitive numerical analysis capabilities. The datasets span diverse domains including healthcare, finance, and industrial applications, with detailed anomaly detection results presented in Table 7 in the Appendix.
4.1 Evaluations
Question-Answering
To evaluate the time series reasoning and question-answering capabilities of TsLLM, we assess its performance using the electrocardiogram question-answering (ECG-QA) benchmark (Oh et al., 2023) comprised of ECG recordings, questions, and answers across several question types: yes/no verification (single-verify), binary choice selection (single-choose), and open-ended queries (single-query). We compared TsLLM performance (zero-shot and few-shot) based on exact match accuracy to top-performing specialized models trained on the ECG-QA dataset, such as M3AE (Chen et al., 2022), MedViLL (Oh et al., 2023), SE-WRN + GPT-4 (Oh et al., 2023), Q-Heart (Pham et al., 2025), and ECG-LM (Yang et al., 2025). All these models are explicitly trained for ECG-QA, whereas TsLLM is exposed to ECG-QA during training, but only a small fraction (4%) of its training tokens originated from this dataset. The performance of TsLLM is further enhanced through in-context few-shot learning using just four examples from the same question type. As Table 1 illustrates, despite the lack of fine-tuning, the models perform well in the category, and TsLLM (few-shot) achieves the best or second-best performance across all queries. While the table focuses on a subset of the question types from this dataset, TsLLM can be applied to other question types with the proper prompt format, demonstrating the ability of the models to handle complex tasks that require expertise.
Table 1: QA accuracy comparison on ECG-QA. The best performance is in bold and the second best is underlined. TsLLM, particularly with few-shot prompting, achieves high performance despite not being fine-tuned for ECG-QA.
Question M3AE MedViLL SE-WRN Q-Heart ECG-LM TsLLM TsLLM
+ GPT-4 (few-shot)
S-Verify 74.6 73.9 71.0 90.9 75.8 77.1 85.3
S-Choose 57.1 54.1 48.1 62.6 57.4 58.9 60.8
S-Query 41.0 40.4 35.7 32.9 39.9 48.2 49.0
Average 57.6 56.1 51.6 61.4 57.7 61.4 65.0
We also evaluate TsLLM on TimeSeriesExam (Cai et al., 2024), a multiple-choice question answering benchmark that tests LLMs’ time series understanding across diverse settings. TsLLM performs far better than even frontier models many times larger than it, demonstrating the effectiveness of our encoding and training approach.
Table 2: Accuracy on TimeSeriesExam.
Accuracy
Model Easy Hard
Gemma 3 27B 0.410 0.368
Gemini 2.5 Flash 0.333 0.237
o4-mini 0.421 0.596
o3 0.447 0.526
TsLLM 0.873 0.855
Classification
For classification tasks, we evaluate our model on the UCR-TSC dataset. Performance is assessed using the mean and median accuracy across sub-datasets, ensuring robustness. Results can be found in Table 3.
Table 3: Time series classification accuracy on the UCR-TSC benchmark collection. TsLLM achieves competitive performance (81.2% mean accuracy with few-shot prompting) without task-specific architectures, demonstrating effective transfer of language model capabilities to temporal pattern recognition.
Metric DTW CNN T-Loss TS2Vec GPT4TS Times- MOM- TsLLM TsLLM
Net ENT0 (1S) (FS)
Mean 0.764 0.751 0.833 0.851 0.566 0.572 0.794 0.792 0.812
Median 0.768 0.773 0.849 0.871 0.583 0.565 0.815 0.798 0.823
Forecasting
We evaluate forecasting performance on the GIFT-Eval dataset. Performance is measured using a relative Mean Absolute Percent Error (rMAPE) normalized against the seasonal naive baseline. All baseline results are as reported in Aksu et al. (2024). We further break down performance by the number of variates (univariate vs multivariate) and prediction length, as illustrated in Table 4, where TsLLM performs competitively with other models although not completely outperforming all.
Table 4: Forecasting performance on GIFT-Eval benchmark across different data characteristics. Values show relative MAPE (lower is better) normalized to seasonal naive baseline, following (Aksu et al., 2024). TsLLM with few-shot prompting (FS) achieves best-in-class performance on univariate series (0.78) while remaining competitive overall (0.87), with particularly strong results on long-horizon forecasts.
Subset
Naive
Seasonal Naive
TFT
PatchTST
iTransformer
TimesFM
VisionTS
Chronos Base
Chronos Large
Moirai Base
Moirai Large
TsLLM
TsLLM (F-S)
GIFT-Eval
All 1.20 1.00 1.12 0.86 0.98 1.25 0.98 0.94 0.93 1.01 0.86 0.89 0.87
Univariate 1.19 1.00 0.89 0.85 0.85 0.87 0.93 0.81 0.80 0.95 0.81 0.82 0.78
Multivariate 1.21 1.00 1.40 0.88 1.15 1.71 1.04 1.11 1.10 1.09 0.93 0.97 0.99
L Horizon 1.35 1.00 1.16 0.96 0.99 1.80 1.03 0.99 1.00 1.09 1.01 1.00 0.96
M Horizon 1.30 1.00 1.26 0.90 1.02 1.52 1.02 1.38 1.33 1.11 0.96 0.96 0.93
S Horizon 1.10 1.00 1.05 0.80 0.97 0.93 0.95 0.75 0.75 0.94 0.77 0.83 0.82
ETTh1 48 1.62 1.00 0.94 0.78 0.81 0.84 0.93 0.76 0.83 0.95 0.87 0.90 0.83
M4 hourly 48 2.42 1.00 1.87 1.07 1.13 0.65 0.81 0.54 0.54 0.82 0.68 0.74 0.65
Electricity 480 1.88 1.00 0.89 0.75 0.84 1.15 0.79 0.82 0.82 0.94 0.79 0.79 0.73
Electricity 720 2.58 1.00 0.77 0.75 0.91 1.25 0.92 0.87 0.87 0.95 0.86 0.78 0.74
4.2 Ablation Studies
Table 5: Forecast performance on GIFT-Eval for each experiment in our ablation study, as measured by relative MAPE.
Ablation Value rMAPE Relative Change
Baseline 0.94
LLM SmolLM2-135M 0.99 -5.5%
LLM SmolLM2-1.7B 0.92 +2.1%
LLM Llama 3.2 3B 0.89 +5.3%
Adapter Linear 0.94 -0.2%
Adapter Cross Attention 0.96 -2.4%
Encoder VQ-VAE 1.24 -31.9%
In order to analyze the performance characteristics of different aspects of the TsLLM architecture and training procedure, we conduct ablation studies on our methodology at a smaller scale. In Table 5, we report the results of our primary findings. The baseline for this experiment uses SmolLM2-360M as the pretrained LLM backbone with an MLP-based adapter for the time series encoder and decoder models, and we utilize the same encoder-decoder pair as in previous experiments. These results demonstrate that the size and performance of the backbone LLM matter significantly, suggesting that TsLLM can be scaled up further to improve performance. They also show that a single linear layer performs nearly as well as the MLP adapter, although we have seen the MLP improves performance on other tasks. Finally, using an encoder-decoder pair with quantized latent embeddings currently has a big negative impact on performance, although they are commonly used in vision-language models. This likely could be improved in future work, but highlights the importance of precise, high-fidelity encoder and decoder models for time series data.
Time Series Encoder-Decoder Training
An ablation study on time series encoder-decoder reconstruction performance (Table 6) reveals a critical trade-off between the reconstruction performance of the encoder-decoder pair and downstream performance on practical tasks. We look at the reconstruction MSE on our VAE validation data across encoder-decoder pairs with different hyperparameter settings, as well as the accuracy of small-scale implementations of TsLLM trained using each of these pairs. This table makes clear that it is possible to improve the reconstruction ability of the VAE significantly by reducing the patch size or $\beta$, but this leads to more granular representations of the time series that are not as informative or effective when integrated into TsLLM. The representation abstraction can also be increased by increasing $\beta$ or selecting a larger patch size, but this hurts reconstruction, which also hurts performance of TsLLM. We have optimized for a balance between these objectives in order to maximize performance of TsLLM on real-world tasks.
Table 6: Trade-off between reconstruction quality and downstream task performance for different encoder-decoder configurations. Lower reconstruction MSE does not guarantee better ECG-QA accuracy – our $\beta$-VAE baseline achieves optimal task performance despite higher reconstruction error than a standard autoencoder. Color coding: green indicates improvement, red indicates degradation relative to baseline.
Ablation Baseline Patch size $\beta$ Objective # of layers
Value 16 64 1.0 1.8 AE VQ-VAE 12 48
Recon. MSE 0.217 0.151 0.308 0.173 0.367 0.005 0.432 0.217 0.216
% Change -30% +42% -20% +69% -97% 99% 0% -1%
ECG-QA Acc. 55.9 54.1 53.5 55.5 55.3 54.0 51.8 55.6 54.9
% Change -3.2% -4.3% -0.7% -1.1% -3.4% -7.3% -0.5% -1.8%
5 Discussion
Training LLMs to understand and analyze numerical time series data represents a paradigm shift in time series analysis, moving beyond traditional approaches that treat numerical data in isolation. By augmenting an LLM with time series perception capabilities, we enable new analytical workflows at the intersection of temporal reasoning and natural language understanding. This approach parallels the successful development of vision-language models, which have proven valuable not by achieving state-of-the-art performance on traditional computer vision benchmarks, but by enabling entirely new capabilities such as visual question answering, image captioning, and multimodal reasoning. Similarly, TsLLM’s value lies not in marginal improvements on forecasting accuracy or classification F1 scores, but in its ability to answer questions about temporal patterns, incorporate unstructured contextual information into predictions, and provide natural language explanations for time series phenomena.
The implications of this paradigm shift extend beyond technical capabilities. TsLLM democratizes time series analysis by allowing domain experts – clinicians, financial analysts, environmental scientists – to directly query and understand temporal data through natural language, without requiring specialized data science training. A cardiologist can ask “What abnormalities are present in this ECG, and are they consistent with the patient’s medication history?” A supply chain analyst can request “Forecast next month’s demand considering recent customer reviews and competitor actions.” These interactions were previously impossible with traditional time series models, which require carefully structured numerical inputs and produce only numerical outputs.
Our approach is fundamentally complementary to, rather than competitive with, specialized time series models. Just as vision-language models have not replaced CNNs for image classification but have opened new application domains, TsLLM does not aim to replace ARIMA models or neural forecasters for pure numerical prediction tasks. Instead, it addresses a different class of problems: those requiring the integration of temporal analysis with contextual understanding, domain knowledge, and natural language interaction. This positioning explains why we do not emphasize traditional benchmark performance as our primary evaluation metric – excelling at these benchmarks is not our goal, just as vision-language models do not target ImageNet state-of-the-art as their primary objective.
Limitations
Several limitations of TsLLM warrant discussion. First, TsLLM requires substantially more computational resources than traditional time series models, both for training and inference. This is an expected tradeoff when leveraging the capabilities of large language models – similar to how vision-language models require vastly more training than specialized image classifiers. The benefit is that TsLLM can be trained once and applied to diverse tasks through in-context learning, amortizing costs across applications and eliminating the need for task-specific model development. Second, while TsLLM demonstrates competitive performance on traditional benchmarks, it does not achieve state-of-the-art results on pure numerical tasks like forecasting or anomaly detection. As previous work has noted (Tan et al., 2024; Merrill et al., 2024), transformers, and LLMs in particular, are not ideally suited for numerical time series data, and we have not overcome this limitation. However, our goal is not to compete with specialized models optimized for specific metrics, but to enable capabilities that those models cannot provide. The performance gap on traditional benchmarks is acceptable given TsLLM’s unique ability to incorporate contextual information, go beyond the standard time series tasks, and answer questions. Our current approach to multivariate time series also has limitations. While effective for independent signals, the model may not fully capture complex interrelationships among variables. This reflects our design choice to handle multivariate complexity within the LLM rather than the encoder, prioritizing flexibility over specialized architectural optimizations. Data availability remains a challenge. High-quality datasets combining time series with natural language descriptions are scarce, requiring us to create our own collection through synthesis and curation. While we have assembled the largest publicly available dataset of interleaved time series and text, it represents only a fraction of potential applications. Expanding coverage to new domains requires substantial effort. Finally, evaluation methodology for time series-language tasks lacks standardization. Unlike pure time series or pure language tasks, there are no established benchmarks for assessing models that bridge both modalities. Existing attempts often rely on synthetic data with questionable quality. Developing rigorous, expert-validated benchmarks for tasks like time series question-answering and contextual forecasting should be a priority for the field.
6 Conclusion
This work presents TsLLM, a large language model augmented with the ability to understand and generate numerical time series signals. By introducing a lightweight patch-based encoder-decoder pair and adapter layers, we enable a pretrained LLM to process interleaved numerical signals and text without causing major degradation to its existing capabilities. A unified autoregressive training objective allows TsLLM to tackle forecasting, classification, anomaly detection and question answering in a single framework, and extensive experiments demonstrate competitive performance on traditional numerical benchmarks while unlocking tasks that require joint temporal and linguistic reasoning.
Our findings highlight three main take-aways. First, high-fidelity numerical encodings are essential: representations that excel at reconstruction translate directly into better downstream performance. Second, the quality of the backbone LLM remains a dominant factor; stronger language models amplify the benefits of specialised perception modules. Third, few-shot prompting can be a practical substitute for task-specific fine-tuning, suggesting that future work can focus on data quality rather than ever larger supervised corpora.
Despite these advances, TsLLM is not a replacement for purpose-built time series models. It lags behind state-of-the-art forecasting models on long-horizon multivariate benchmarks and is much more computationally expensive. We therefore view TsLLM as complementary: it excels when time series data must be interpreted in the context of domain knowledge or when natural-language interaction is required.
Looking ahead, we believe the most fruitful directions lie in (i) better handling of multivariate dependencies, (ii) tighter integration of structured domain knowledge, and (iii) community benchmarks that evaluate combined numerical-linguistic competence. We release code, models and data to facilitate such work and hope that TsLLM serves as a stepping stone toward more capable, transparent and deployable time series foundation models.
Appendix A Implementation Details
We implemented TsLLM using PyTorch and Transformers, with FlashAttention-2 and fully sharded data parallelism for efficient distributed training. The complete training process consumed approximately 30 billion tokens across all stages, requiring approximately 1000 GPU-hours on NVIDIA H100s. Our primary experiments use Qwen2.5-7B [Wang et al., 2024] as the base LLM, chosen for its strong performance on standard benchmarks. We also validated our approach with smaller models and different architectures (SmolLM2-360M [Allal et al., 2025], Llama-3.2-3B [Grattafiori et al., 2024]), before scaling up to this size.
The encoder employed a patch-transformer architecture with 24 layers, an embedding size of 64, and a patch size of 32. Time series inputs were standardized using Z-score normalization. The VAE’s latent space is 16-dimensional, and the KL divergence loss was scaled with a peak $\beta$ coefficient of 1.5. VAE training was conducted for 100 million samples using a batch size of 2048. The optimizer was AdamW, with a learning rate of 5e-4 and a linear warmup scheduler. The mean squared error (MSE) loss was used for reconstruction.
TsLLM was optimized using AdamW, with a learning rate of 1e-5 and gradient norm clipping set to 1.0. The total training dataset comprised 20 billion tokens, with a sequence length of 4096 and a micro-batch size of 4. MSE loss was used for time series next patch prediction on time series tokens, and cross entropy loss was used on all non-padding tokens.
Appendix B Additional Results
In this section we present additional empirical results as a continuation from Section 4. Figure 3 plots the relationship between the patch size set for the time series encoder and decoder, and the resulting reconstruction error (MSE) of the VAE model on a subset of the LOTSA dataset. It demonstrates that for a fixed representation size, patch size and reconstruction performance are inversely correlated, so smaller patch sizes have inherent benefits. However, they also increase the sequence length of a time series, shifting complexity to the LLM, which can harm downstream performance. Finding the right balance between these factors is critical for optimizing capabilities across a range of tasks and datasets.
Figure 3: The relationship between patch size for the time series encoder and reconstruction MSE for the VAE on the validation set.
We also evaluate TsLLM on anomaly detection, specifically on the MSL dataset [Hundman et al., 2018], as shown in Table 7. While our approach achieves a competitive F1 score of 73.3, there remains a performance gap compared to state-of-the-art models like UniTS-FT (81.2) and iTransformer-FT (80.4). This suggests that while TsLLM’s general-purpose architecture provides reasonable anomaly detection capabilities, there may be room for incorporating specialized mechanisms to better identify temporal anomalies. The results demonstrate TsLLM’s versatility across different time series tasks, though specialized models may still hold advantages in specific domains.
Table 7: Anomaly detection performance on MSL.
Model MSL (F1 score)
Anomaly Trans. 78.0
TimesNet-FT 33.9
iTransformer-FT 80.4
PatchTST-FT 79.9
UniTS-PMT 75.4
UniTS-FT 81.2
TsLLM 73.3
Appendix C Datasets
Our training approach leverages a comprehensive collection of both curated and synthetic datasets, spanning multiple domains and time series analysis tasks. We strategically combine established benchmarks with novel synthetic data generation techniques to ensure broad coverage of time series reasoning capabilities while maintaining data quality and diversity. This section provides detailed descriptions of our data sources, organized by modality and generation approach.
C.1 Training Data Overview
Table 8 summarizes the key statistics for all datasets used in training TsLLM. Our training corpus encompasses over 8 billion total tokens across 20 distinct datasets, spanning diverse domains including medical signals, financial data, environmental measurements, and industrial sensors. The datasets support multiple task types including forecasting, classification, anomaly detection, segmentation, and question-answering, enabling TsLLM to develop comprehensive time series analysis capabilities.
Name Domain Tasks Total Tokens Response Tokens Text/TS Tokens Samples
ECG-QA ECG Q-A 110,278,285 607,619 (34,289,034/75,381,632) 159,306
GIFT Multi-Domain Forecasting 10,380,569 7,724,958 (0/2,655,611) 96,307
LOTSA Multi-Domain Forecasting 7,951,924,108 5,961,122,236 (0/1,990,801,872) 74,805,850
News-Stocks Finance Forecasting 124,179,199 3,464,920 (118,277,891/2,436,388) 61,922
PTBXL-Classific. ECG Classification 20,615,930 296,710 (3,577,588/16,741,632) 21,799
PTBXL-Forecast. ECG Forecasting 22,482,365 5,740,733 (0/16,741,632) 21,799
PTBXL-Rep. Gen. ECG Report Gen. 23,155,582 3,689,075 (2,724,875/16,741,632) 21,799
GIFT-Eval (Val) Multi-Domain Forecasting 39,338,366 27,820,967 (0/11,517,399) 346,955
TTC-Medical Medical Mixed 628,367 232,387 (390,936/5,044) 330
TTC-Climate Climate Mixed 2,132,627 559,584 (1,563,351/9,692) 1,320
UCR-TSC Time Series Classification 68,966,571 797,925 (50,634,942/17,533,704) 60,555
UTSD Multi-Domain Forecasting 78,164,051 51,392,965 (0/26,771,086) 873,621
UEA Multi-Domain Classification 29,306,232 1,821,291 (16,888,337/10,596,604) 22,606
Child Sleep Detct. Sensor Segmentation 916,281 52,061 (531,027/333,193) 5,331
PSM Machine Anomaly 704,856 240,546 (285,299/179,011) 220,322
MSL Spacecraft Anomaly 893,631 304,969 (361,708/226,954) 132,046
SMAP Spacecraft Anomaly 1,731,267 590,829 (700,751/439,687) 562,800
SWaT Infrastructure Anomaly 6,106,917 2,084,106 (2,471,848/1,550,963) 944,920
SMD Machine Anomaly 6,624,762 2,260,831 (2,681,452/1,682,479) 1,416,825
Synthetic (Ours) Multi-Domain Mixed 7,481,857,697 7,011,433,500 (368,202,956/102,221,241) 6,273,306
Table 8: Comprehensive statistics for all datasets used in TsLLM training, showing the distribution of tokens across text and time series modalities, along with task diversity across domains.
C.2 Unimodal Time Series Data
Our approach leverages a diverse collection of unimodal time series datasets spanning multiple domains and analytical tasks. These datasets provide the foundation for training TsLLM’s time series understanding capabilities across both unlabeled pretraining data and task-specific labeled benchmarks.
Large-Scale Pretraining Data
For large-scale pretraining, we utilize the GIFT-Eval benchmark [Aksu et al., 2024], a comprehensive collection of 144,000 time series across 23 datasets spanning seven domains. GIFT-Eval provides dedicated splits for pretraining, training, and testing, making it particularly suitable for developing and evaluating general-purpose time series models. The benchmark’s temporal granularity ranges from secondly to yearly frequencies, incorporating both univariate and multivariate sequences with forecast horizons varying from short-term to long-term predictions. With over 230 billion data points, this dataset provides the scale and diversity necessary for training foundation models. We use it extensively for pre-training our time series encoder and decoder models, as well as training TsLLM for forecasting tasks, while carefully ensuring no data leakage into the GIFT-Eval test set used in our evaluation (Section 4).
We also leverage two additional large-scale datasets for comprehensive coverage: the Unified Time Series Dataset (UTSD) from Tsinghua University’s THUML lab, a large-scale collection spanning seven domains (Energy, Environment, Health, IoT, Nature, Transportation, and Web) with up to one billion time points that combines publicly available online data with empirical measurements from machine operations, and LOTSA from Salesforce, which encompasses approximately 27 million observations across nine domains (Energy, Transport, Climate, CloudOps, Web, Sales, Nature, Economics/Finance, and Healthcare) with sampling frequencies ranging from seconds to years.
Classification Benchmarks
For time series classification tasks, we employ two principal benchmark collections that have become standard in the field. The UCR Time Series Classification Archive [Dau et al., 2018] comprises 128 datasets spanning diverse applications including electrocardiograms, motion capture sequences, sensor readings, and spectrographs. These datasets vary significantly in length, class distribution, and domain-specific characteristics, providing comprehensive coverage for univariate time series classification. Complementing this, the UEA Time Series Classification Archive [Bagnall et al., 2018] extends our capabilities to multivariate time series, offering 30 datasets that capture complex temporal patterns in applications such as human activity recognition, motion capture, and handwriting analysis. Together, these repositories represent the gold standard for time series classification research, with the UCR archive focusing on univariate sequences and the UEA archive addressing more complex real-world scenarios where multiple variables are tracked concurrently.
Specialized Task Datasets
For segmentation tasks, we employ the Child Mind Institute’s "Detect Sleep States" dataset [Esper et al., 2023], which consists of approximately 500 multi-day accelerometer recordings captured from wrist-worn devices, with each series corresponding to a unique participant. The dataset provides expert annotations for sleep onset and wakeup events, enabling comprehensive evaluation of temporal segmentation capabilities in physiological monitoring applications.
For anomaly detection, we incorporate several benchmark datasets representing different operational contexts and domains. These include NASA’s Mars Science Laboratory (MSL) and Soil Moisture Active Passive (SMAP) telemetry datasets [Hundman et al., 2018], which provide spacecraft sensor data with labeled anomalies; eBay’s Pooled Server Metrics (PSM) dataset [Abdulaal et al., 2021], containing server performance metrics; the Server Machine Dataset (SMD) [Su et al., 2019], which includes machine operation data; and the Secure Water Treatment (SWaT) dataset [Mathur and Tippenhauer, 2016], representing critical infrastructure monitoring. This diverse collection ensures robust anomaly detection capabilities across industrial, space, and cybersecurity applications.
Medical Time Series Data
We also utilize the PTB-XL dataset [Wagner et al., 2022, Goldberger et al., 2000], which consists of 21,799 clinical 12-lead ECGs collected from 18,869 patients. Each recording spans 10 seconds and includes expert annotations from up to two cardiologists, using 71 distinct SCP-ECG standard statements. The dataset includes comprehensive metadata including patient demographics and diagnostic information, making it particularly valuable for developing medical time series analysis capabilities. From this dataset, we developed specialized variants optimized for different tasks: one for time series classification tasks and another for time series forecasting applications, as detailed in our data synthesis section.
Text-Only Training Data
To ensure TsLLM maintains strong text generation capabilities alongside its time series analysis skills, we incorporate two standard text datasets into our training mixture. We utilize a subset of FineWeb-Edu [Lozhkov et al., 2024], a large-scale educational text corpus containing high-quality educational content curated from various web sources and specifically filtered for instructional content, explanations, and academic materials. Additionally, we include SmolTalk [Allal et al., 2024], a synthetic dataset of 1M instruction-following examples created for supervised fine-tuning, covering tasks like summarization and reasoning. This combination ensures that our model’s natural language capabilities remain robust while learning time series-specific skills.
C.3 Paired Time Series and Text Data
Multimodal datasets that combine time series with textual information are essential for training TsLLM’s ability to reason about temporal patterns in context. Our approach leverages both existing paired datasets and novel synthetic generation techniques to create comprehensive training data for time series-text understanding.
Existing Multimodal Datasets
Existing datasets that pair time series with textual data have been instrumental in developing models that can bridge numerical and linguistic understanding. The ECG-QA dataset [Oh et al., 2023] provides a specialized medical dataset that pairs electrocardiogram readings with natural language questions and answers about cardiac conditions and diagnoses. Built from PTB-XL [Wagner et al., 2022] and MIMIC-IV-ECG [Gow et al., 2023] databases, it includes both templated and naturally-phrased questions that test understanding of ECG features like heart rhythm patterns.
The TextTime Corpus (TTC) [Kim et al., 2024] offers broader domain coverage by pairing two types of temporal data: (1) daily weather measurements with forecast discussions and (2) patient vital signs paired with clinical notes from MIMIC-III [Johnson et al., 2016]. This diversity in both domains and text types provides valuable training examples for contextual reasoning about time series data.
Time-MMD [Liu et al., 2024c] extends multimodal capabilities by combining numerical and textual sequences across 9 domains, incorporating corresponding reports and search results that provide rich contextual information for time series analysis.
Financial Dataset Development
For our data collection efforts, we focused particularly on the financial domain, creating a novel dataset that combines S&P 500 daily closing prices with corresponding news articles. This pairing enables exploration of contextual forecasting tasks where models must consider both numerical trends and relevant textual information when making predictions.
Our financial dataset construction process involves several key steps: first, we collect historical S&P 500 price data; second, we gather contemporaneous news articles that might impact market movements; and finally, we align these data sources temporally to create paired examples suitable for contextual forecasting tasks.
We constructed this dataset using news data from two sources with different granularities: The Bloomberg dataset [Genloop, 2024] provides full articles with multiple stock symbol associations, while the Benzinga dataset [Aenlle, 2020] contributes headlines with single stock symbol associations. The dataset includes daily closing prices and company fundamentals (sector, industry, country, and business descriptions). Each forecasting task combines historical price data and news context (up to 4 Bloomberg articles or 16 Benzinga headlines) to predict a 32-day forward price sequence. We employ a minimum context window of 32 days with prediction windows spaced at 16-day intervals, ensuring all contextual information precedes the target sequence.
C.4 Synthetic Data Generation
To ensure comprehensive coverage of time series analysis tasks and provide sufficient training data for multimodal understanding, we developed a systematic data synthesis approach encompassing three main components: augmentation of existing time series datasets, foundational text generation tasks, and large-scale question-answering synthesis.
PTB-XL Dataset Augmentation
We created three synthetic variants of the PTB-XL dataset to expand its applicability beyond traditional classification tasks. Using Llama 3.3 70B Instruct [Grattafiori et al., 2024] with controlled generation parameters (temperature=0.7, top-k=40), we systematically generated textual content paired with the existing ECG recordings.
The first variant, designed for contextual forecasting, augments each ECG with multiple generated reports using different templates including basic summaries, highlight-focused reports, and interpretative reports with clinical context. Each template was designed to handle both English and German inputs while maintaining English output consistency. The second variant, optimized for classification tasks, pairs each ECG with demographic information (age, sex, height, weight) while excluding statement fields and signal-derived features. This cleaned subset ensures models must rely primarily on raw signal patterns rather than pre-computed features, providing a more challenging and realistic evaluation scenario.
The third variant supports report generation tasks by combining ECG signals with reports generated under specific constraints. These include structured responses containing only signal-derived information, narrative reports presenting findings as flowing clinical stories, and interpretative reports providing broader clinical context. Each record includes both the generation prompt and the resulting report, enabling comprehensive evaluation of medical time series understanding and generation capabilities.
Foundational Time Series Tasks
To establish and evaluate basic time series manipulation capabilities, we created a synthetic instruction-following dataset using time series sampled from the GIFT-Eval pretraining split. Employing various state-of-the-art LLMs including Gemini 1.5 Pro [Team et al., 2023], we generated diverse instruction-response pairs across four fundamental tasks that serve as building blocks for more complex time series operations.
The first task involves encoding time series as text, where numerical sequences are converted to comma-separated string representations with consistent precision and formatting. The complementary second task requires reconstructing time series from their textual encodings, testing the model’s ability to accurately parse and interpret numerical sequences. The third task focuses on basic statistical computations, specifically calculating fundamental measures like mean and standard deviation of given time series. The fourth task extends this to comprehensive statistical analysis, requiring the model to compute and return a structured JSON object containing seven statistical measures including median, quartiles, and extrema.
This collection of foundational tasks serves as a warm-up phase in training, establishing essential capabilities in numerical manipulation, statistical analysis, and format conversion that underpin more sophisticated time series reasoning and analysis operations.
Large-Scale Question-Answering Synthesis
To create a comprehensive training dataset for time series question-answering capabilities, we synthesized a large-scale dataset comprising over 2 million question-answer pairs, totaling approximately 10 billion tokens. This dataset was generated through a systematic combination of carefully curated question templates, diverse time series data, and responses from state-of-the-art language models. All time series data was sampled from the GIFT-Eval pretrain split, with rigorous filtering applied to ensure data quality and maintain consistent sequence lengths suitable for training.
For time series representation, we implemented two complementary encoding schemes to maximize model flexibility. The text encoding represents time series values as comma-separated numerical sequences with configurable precision and normalization options, accompanied by rich metadata including sampling frequency, domain information, and basic statistical summaries. The image encoding visualizes time series as line plots encoded as RGB images, enabling compatibility with vision-language models and providing an alternative modality for time series understanding.
Our question-generation system employs template-based sampling with carefully weighted instruction types and styles. The templates comprehensively cover various analytical tasks ranging from trend identification and pattern recognition to statistical analysis and anomaly detection. Each generated question is paired with relevant time series metadata and optional statistical summaries to provide sufficient context for accurate and informative responses. This approach ensures broad coverage of different analytical scenarios while maintaining consistency in question quality and complexity.
To generate high-quality responses, we employed a diverse ensemble of state-of-the-art language models including text-only models (Llama 3.3 70B [Grattafiori et al., 2024]), vision-language models (Qwen2-VL-72B [Wang et al., 2024]), and frontier models (Gemini 2.0 Flash [Team et al., 2023]). While synthetic responses may contain some inaccuracies, recent findings [Bansal et al., 2024] suggest that larger quantities of moderate-quality training data can be more effective than smaller sets of perfect examples for LLM training. The resulting scale and diversity of our dataset provides comprehensive coverage of different analytical scenarios, making it particularly suitable for training models in time series analysis and reasoning tasks.
Appendix D Data Samples
The training data consists of textual instructions followed by a time series sequence, marked with a special token. The corresponding labels provide the indices of sleep onset and wake-up points. Example training instances is shown below:
Child Sleep Detection
Input:
"Instruction: Identify sleep onset and wake-up points in the following time series.
Time series signal: <|ts|>"
Label:
"onset indices": [252, 469]
"wakeup indices": [347, 610]
Gift-eval
Input:
<|begin_of_text|> Time series metadata: dataset = Monash M3 Quarterly,
start = 1977-01-01T00:00:00, end = 1993-07-01T00:00:00,
sample frequency = Q-DEC, length = 66 points
Time series signal: <|ts|> <|end_of_text|>
Synthetic QA
Input:
<|begin_of_text|> You are a helpful AI assistant trained to understand numerical time series data.
Prompt: Describe how this time series might continue if extended.
Series metadata:
domain: Nature
dataset: Star Light Curves
Time series:
Time series: <|ts|>
Response:
The time series you’ve provided appears to represent the brightness (or magnitude) of a star over time, specifically from the dataset of star light curves. This kind of data is crucial in astrophysics for understanding the behavior of stars, such as their variability and periodicity.
Analysis of the Time Series:
1. Trend and Pattern: The series starts with a gradual decrease in brightness, reaching a minimum value, and then gradually increasing back to a higher brightness. This suggests a periodic or cyclic pattern, which is common in variable stars.
2. Periodicity: The pattern of decrease and increase repeats, indicating that the star might be a pulsating variable star (like a Cepheid variable or RR Lyrae star) or a star with an eclipsing companion (like an eclipsing binary system).
3. Amplitude: The amplitude of the variation is significant, ranging from approximately -0.57 to 2.80 in the magnitude scale. This is typical for stars with strong variability.
4. Duration of Cycle: The complete cycle from peak to trough and back to peak appears to be quite long, suggesting a long-period variable star.
Prediction of Future Behavior:
Given the periodic nature of the time series, we can predict that the pattern will likely continue in a similar fashion:
1. Continuation of the Cycle: The series will likely continue to show the same pattern of gradual increase and decrease in brightness. After reaching the latest minimum value, it will start to increase again, reaching another peak, and then decrease to another trough.
2. Amplitude and Frequency: The amplitude of the variations will likely remain consistent with the observed pattern, with the brightness increasing to a peak similar to the previous maximum (around 2.80) and then decreasing to a similar minimum (around -0.57).
3. Possible Variations: While the overall pattern is expected to continue, there may be small variations or irregularities due to factors such as:
- Noise: Random fluctuations that can occur due to observational errors or intrinsic variability.
- Long-term Trends: Gradual changes in the star’s behavior over time, which might cause the amplitude or period to change slightly.
- External Factors: Events like stellar flares or interactions with a companion star could introduce anomalies.
Visualization:
To better understand the continuation, a visual representation of the time series would show a wave-like pattern, with the latest part of the series starting to increase in brightness after reaching a minimum. The next cycle would follow a similar trajectory, peaking at a similar value and then decreasing again.
Conclusion:
The time series is likely to continue with a periodic pattern of increasing and decreasing brightness, maintaining a similar amplitude and period as observed. Any deviations from this pattern would be interesting and could indicate changes in the star’s behavior, which would be valuable for further astrophysical study.
<|end_of_text|>
Appendix E Broader Impacts
Our work introduces a novel architecture leveraging established open-source model components to achieve new capabilities in processing and generating mixed-modal time series and text. While we build upon models that have undergone community vetting, our architecture’s application to diverse domains, including healthcare and finance, warrants a careful consideration of broader impacts.
Positively, this approach offers significant potential. By integrating textual context with time series analysis, it can enhance the interpretability and utility of complex data, for example, by aiding in the summarization of ECG signals or providing textual explanations for environmental data trends. This could improve decision-making and accessibility across scientific and industrial applications. The ability to generate multimodal data also holds promise for data augmentation and system testing.
However, we recognize that the combination of text and time series generation, even using existing components, presents specific risks. Our primary concern is the risk of halucinations if this line of work is used for safety-critical problems. Progress on significantly reducing hallucinations in LLMs, as well as the development of robust safeguards, are important areas for future work, and necessary before this line of work can be used in safety-critical domains. Furthermore, as with any model trained on public data, there’s a risk of perpetuating inherent biases if not carefully managed, particularly in sensitive applications like healthcare. Errors in critical tasks, though a risk in any system, could also have significant consequences.
Our research focuses on the architectural innovation, and we are not releasing a new frontier model. Nevertheless, we believe any future deployment of systems based on these principles should incorporate robust safeguards. These would include careful dataset curation to mitigate bias, human-in-the-loop systems for critical decision-making (especially in healthcare), and ongoing research into detecting AI-generated multimodal content. We encourage responsible development practices as this line of research progresses.
From an ethical and societal perspective, TsLLM can democratize powerful modeling tools across healthcare, environmental sciences, and industry, especially in resource-limited settings. However, risks include potential misuse and overreliance in critical decision-making. Future work should improve interpretability, robustness to distributional shifts, and domain-specific validation to ensure responsible use.